OpenStudy (anonymous):

what is the graph of 5x=2y-20?

OpenStudy (anonymous):

Rearranging it into a y=mx+b form might help you visualize it

OpenStudy (anonymous):

how

OpenStudy (anonymous):

5x=2y-20 2y=5x+20 y=2.5x+10